# Project Page: https://subjectdrive.github.io/

### v1

@@ Abstract
Keywords:
Video Generation Autonomous Driving Data Scaling Subject Control
Figure 1: An overview of the proposed SubjectDrive framework and its effectiveness in enhancing BEV perception tasks. (a) Traditional data generation framework that uses the control sequence and sampling noise to generate synthetic data. (b) Compared with the traditional framework, our SubjectDrive introduces additional synthesis diversity by incorporating extra subject control. (c)-(d) Evaluation of detection and tracking performance with data scaling. (e) Illustration of using the SubjectDrive framework to produce perception training data in autonomous driving.
Equal contribution. This work was done during the internship at MEGVII Technology. Corresponding author.
1 Introduction
@@
Deep generative models [34, 4, 47, 48, 37, 15] have made significant progress recently, excelling in producing high-quality, realistic visual contents. Diffusion models [37, 15], a key contributor to this advancement, are noted for their stable and superior-quality sample generation. Leveraging modern diffusion models, recent breakthroughs in controllable technology [48] now enable precise, flexible content customization. These advancements make it possible to create synthetic samples nearly indistinguishable from human-annotated data, sparking increased interest in using generative models for effective natural data synthesis in complex discriminative tasks [36, 2, 30].
BEV (Bird’s Eye View) perception in autonomous driving is one of the most crucial yet challenging discriminative tasks in computer vision. Training robust BEV perception models, such as those for 3D object detection [40, 32, 23, 24, 25] and tracking [32, 9, 29], demands large-scale annotated samples. However, acquiring such datasets is not only costly and time-intensive but also encumbered by significant concerns regarding data privacy and usage rights. These barriers significantly hinder the processes of data collection, labeling, release, and exchange, ultimately slowing progress in the field. Amidst the significant strides in deep generative models, a pressing question emerges: Could these innovations transform the data production pipelines for autonomous driving, thereby catalyzing its advancement?
There have been a few initial attempts at exploring generative data in autonomous driving. Early studies focused on creating street-view images to enhance image-based BEV perception methods, introducing solutions such as BEVGen [38], BEVControl [44], and MagicDrive [11]. More recent efforts have ventured into generating driving scene videos [43, 20, 27], aiming to bolster more sophisticated temporal BEV perception models [40, 23]. A notable example, Panacea [43], excels in producing multi-view videos with BEV controlling capabilities, significantly aiding in the training of the advanced BEV perception model [40].
Albeit promising, most existing works focus on small-scale training [43, 20, 27], which fails to fully exploit the potential of generative data in producing an inexhaustible supply of samples. Notably, early methods have yet to demonstrate superior performance in supporting downstream perception tasks, especially when compared to models pre-trained on large-scale, realistic datasets. For instance, the advanced generative method Panacea achieves a nuScenes Detection Score (NDS) of 49.2 in the 3D detection tasks. This score is lower than that of the model pre-trained on the nuImages [5] dataset, as illustrated in Fig. 1(c).
@@
In this paper, we tackle the challenge of scaling generative data for autonomous driving. We present SubjectDrive, an advanced video generation framework designed to enhance the scalability of generative models. Our initial findings reveal that conventional video generation pipelines struggle to scale effectively with increased data volumes. For instance, as shown in Fig. 1(c-d), tripling the generative data via Panacea yields only a minor improvement. To overcome this limitation, we propose a novel generation framework centered on augmenting sampling diversity. Specifically, we integrate a feature termed subject control into existing generation pipelines, as illustrated in Fig. 1(b). This feature empowers generative models to manipulate the diversity of the synthesis process by providing a mechanism to dictate the visual appearance of foreground elements in generated samples. This innovative feature enables the blending of the inherent stochastic nature of the generative process with the diversity drawn from external data sources, thereby crafting a more powerful model capable of producing scalable and diverse samples.
@@
Specifically, SubjectDrive is designed around a trio of innovative modules that collectively enable robust subject control capabilities. Initially, the model leverages a subject prompt adapter, integrating subject control seamlessly with the existing text-conditioned branch. Subsequently, to enhance the model’s ability to inject spatial information, we introduce a subject visual adapter that directly utilizes visual features, incorporating them into the existing diffusion U-Net architecture. Lastly, to ensure consistent injection of these features over time, we deploy augmented temporal attention that expands the model’s temporal-spatial context. Together, these modules empower SubjectDrive to perform subject-conditioned video generation with compelling efficacy, as demonstrated by the visualizations in Fig. 1(e).
@@
We have conducted extensive experiments to validate the effectiveness of our proposed method on the widely used nuScenes [5] dataset. Compared to existing methods, SubjectDrive not only achieves superior performance but also offers improved scalability. Remarkably, our method represents the first generative approach capable of enhancing the performance of perception models beyond what is possible with pre-trained models on the nuImages dataset. These outstanding results underscore the potential of generative data to revolutionize autonomous driving technologies, marking a promising path forward in this field.
2 Related Work
2.1 Scalable Data Synthesis for Autonomous Driving
@@
The use of synthetic data has been widely explored for visual perception tasks that require great of efforts in labeled data collection. Examples include applications in image classification [2, 36], semantic segmentation [30], and visual object tracking [19]. Recently, there has been a burgeoning interest in employing generative data for the challenging BEV perception tasks [40, 23, 49], which demand precise geometric and appearance alignment from generative models. Initial works aim at synthesizing street-view images [38, 44, 11], with subsequent studies extending into the generation of driving scene videos [43, 27, 41]. Despite these successful efforts, the majority of existing research has been limited to producing a small quantity of training samples, not fully exploiting the generative models’ capacity to offer an unlimited reservoir of samples. Data scaling efforts [2] are sparse and, thus far, not particularly successful. It’s imperative to recognize that recent breakthroughs highlight the paramount importance of data scaling in training sophisticated deep learning models, such as GPT [1], SVD [3], and Sora [31]. Therefore, addressing the challenge of amplifying generative data volumes is of essence. This work delves into scalable data synthesis for autonomous driving, providing analysis results and proposing an enhanced framework.
2.2 Diffusion-based Generative Models with Subject Controls
@@
Diffusion models with subject control are designed to embed target subjects into generated visual content, guided by reference images [8]. These approaches [12, 18, 21, 7, 28] initially emerged in image generation research to facilitate identity-preserving applications. Early methods [12, 18, 21, 10, 35] adopted a fine-tuning framework to adjust a specific model capable of generating images of the desired subject. A notable example is DreamBooth [35], which fine-tunes a diffusion model using a small set of subject images. While these fine-tuning methods achieved high-quality results, they were limited by the significant resources required for model tuning, making them impractical for scenarios with numerous target entities. To address these limitations, tuning-free methods [28, 46] were developed. Subject Diffusion [28], for instance, achieved impressive customization by introducing a subject conditioning module and training on a bespoke dataset of subject pairs. Subsequent works, such as Cones2 [26] and others [7], have expanded tuning-free subject control to scenarios involving multiple subjects.
@@
In the video domain, there is also growing interest in subject-controllable generation [16, 6, 42]. For instance, VideoBooth [16] proposed a framework capable of generating consistent videos containing the subjects specified in image prompts. VideoDreamer [6] introduced the Disen-Mix Fine-tuning and Human-in-the-Loop Refine-tuning strategy. Additionally, CustomVideo [42] introduced a multi-subject-driven text-to-video model powered by a simple yet effective co-occurrence and attention control mechanism.
Our work falls into the category of subject-driven video generation. However, unlike previous efforts, we explore the significance of such subject-controlled generation for data scaling, especially in the context of autonomous driving.
3 Method
3.1 Preliminary: Latent Diffusion Models
@@
Diffusion Models (DMs) [37, 15] iteratively denoise a random noise with Gaussion distribution $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ over T steps to generate target data according to Eq. 1, where the functions $\mu_{\theta}$ and $\Sigma_{\theta}$ are derived from the denoising model $\epsilon_{\theta}$. The training of DMs includes both a diffusion process and a denoising process. The diffusion process begins by adding Gaussion noise to the original data sample $x_{0}$ over $t$ steps, controlled by the scheduled noise strength $\beta_{t}$, which can be simplified as Eq. 2 where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\left(1-\beta_{t}\right)$. To learn the denoising process, the denoising model $\mathbf{\epsilon}_{\theta}$, which aims to estimate the original noise $\mathbf{\epsilon}$ from the noisy data $x_{t}$, is optimized by minimizing the loss function as shown in Eq. 3.
@@
Given the significant computational burden of diffusion models in generating high-resolution images or videos, Latent Diffusion Models (LDMs) [34] have become popular in recent studies. With a pre-trained auto-encoder [17] to handle the compression and reconstruction between high-dimensional visual data and low-redundancy latent representations, the diffusion model can concentrate exclusively on generation in latent space, significantly reducing computational costs. Consequently, our method employs LDMs for video generation.
3.2 SubjectDrive
SubjectDrive is designed to enhance the scalability of generative data, thereby promoting perception models for autonomous driving applications. Despite the ability of advanced generative methods [43] to produce high-quality driving-scene videos, they narrowly uplifts the performance on downstream perception tasks as illustrated in Fig. 1(c)(d). We believe this is primarily due to the limited diversity of generated foreground elements, which are crucial for autonomous driving. Thus, different from tradition generation pipeline using control sequence to guide global scene generation, we innovatively integrate subject control mechanism into generation process, allowing for the injection of external subjects from extensive open-source data. The integration of controlled subjects not only boosts controllability but also effectively enhances the diversity of generated foreground elements, which shows strong augmentation for autonomous driving applications.
@@ 3.2 SubjectDrive
3.2.1 Overview
Our framework is built on a text-to-video diffusion model, i.e. Panacea [43], which is a strong baseline for multi-view video generation. The overall training framework of SubjectDrive is illustrated in Fig. 2. The diffused noisy input is fed into the trainable diffusion model to generate latent video under the guidance of text, condition layout, and injected subjects. We inherit the guidance of condition layout with ControlNet [48] from Panacea. For subjects guidance, we first extend text prompt and augment the text embedding of subject part with Subject Prompt Adapter. On the other hand, we insert the gated self-attention layer into diffusion model to enhance the location guidance of subjects captured by Subject Visual Adapter. Furthermore, we replace the conventional temporal 1D attention with proposed Augmented Temporal Attention to improve the temporal consistency of injected subjects.
3.2.2 Subject Prompt Adapter
@@
To achieve this, we first extend the scene prompt with injected subject category description. For example, the prompt for the $N_{th}$ frame that contains $M$ subjects is organised as "[Scene Prompt], including [Subject $X_{1}$], [Subject $X_{2}$]…. [Subject $X_{M}$]." [Subject $X_{i}$] here refers to the category word of subject $X_{i}$, e.g., car, bus. The extended prompt is input into the pre-trained CLIP [33] text encoder to obtain the original text embedding $z^{t}\in{\mathbb{R}^{L\times d}}$, where $L$ represents the length of the text embedding and $d$ is the dimension of the text embedding. To further integrate subject attributes for better expressity, we extract the ID identifier of each subject from condition layout to ensure coherent trajectories across frames, and encode it with a learnable codebook and an ID-adapter comprised of MLP to get the ID embedding $z_{i}^{id}\in{\mathbb{R}^{d}}$ for subject $X_{i}$. Similarly, we encode the image of $X_{i}$ by the CLIP image encoder and an image-adapter to obtain the corresponding visual semantic embeddings $z_{i}^{v}\in{\mathbb{R}^{d}}$.
Figure 2: Overview of SubjectDrive. (a) The pipeline of SubjectDrive involves a frozen auto-encoder and a trainable UNet-based diffusion model. (b) Different control signal sources including extended text prompt, condition layout, and subject bank. (c) The Subject Prompt Adaptor which augments the original text embedding of extended prompt with corresponding ID identifier and visual semantic information to enhance the expressity of subjects. (d) The Subject Visual Adapter which injects location-enhanced subject information into visual features cooperated with gated self-attention. (e) The Augmented Temporal Attention integrates conventional temporal 1D attention with decomposed attention on temporal-horizontal (TX) plane and temporal-vertical (TY) plane to effectively capture long-range movements of subjects.
@@
$\hat{z}_{i}^{t}=\mathrm{MLP}\left([z_{i}^{t}+z_{i}^{id},z_{i}^{v}]\right),\ \ i\in\left\{\text{ Index }_{1,2,\cdots,M}\right\}$ (4)
@@
3.2.3 Subject Visual Adapter
@@
Inspired by [22], the SVA first combines subject images and corresponding locations to form location-enhanced subject embeddings, and then injects them into frames with a control signal to guide the generation of subjects at specified location. To obtain the location-enhanced subject embedding $f^{vl}$, the location of each subject, represented by its coordinates $l=[x_{min},y_{min},x_{max},y_{max}]$ in the current frame, is encoded with Fourier embedding, further integrated into visual embedding $f^{v}$ of corresponding subject as
$f^{vl}=\text{MLP}[f^{v},\text{Fourier}(l)]).$ (5)
To inject the location-enhanced subject embedding into generated frame, we employ the gated self-attention layer [22] that locates between each paired self-attention layer and cross-attention layer in the UNet. In the gated self-attention layer, the location-enhanced subject embedding $f^{vl}$ interacts with frame visual tokens $z$ by an attention operation to capture the dependency between subject embedding and visual tokens, followed by a token selection operation $TS$ to preserve only visual tokens. To adaptively adjust the guidance scale of location information over frames, a gating factor is learned which is operated as
$z=z+\tanh(\gamma)\cdot TS(\text{SelfAttn}([z,f^{vl}])),$ (6)
@@
3.2.4 Augmented Temporal Attention
The Augmented Temporal Attention is designed to effectively capture long-range movement of subjects with feasible computation cost.
@@ 3.2.4 Augmented Temporal Attention
$z^{o}=\text{SelfAttn}(z^{TX})+\text{SelfAttn}(z^{TY})+z^{T}+z^{i}.$ (7)
4 Experiment
4.1 Evaluation Datasets and Metrics
@@
Evaluation Metrics. Following Panacea, we employ perceptual model evaluation to validate two key objectives. First, we assess the controllability of the generative model, specifically how closely the data generated aligns with given BEV annotations. We utilize a pre-trained StreamPETR model to compare the validation performance of our generated data with that of real data. Second, we explore the extent to which the generative model’s data improves the performance of the perceptual model. We conduct evaluations on both detection and tracking tasks. For the detection task, we use the nuScenes Detection Score (NDS), mean Average Precision (mAP), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE) as metrics. For the tracking task, our metrics include Average Multi-Object Tracking Accuracy (AMOTA), Average Multi-Object Tracking Precision (AMOTP), Recall (RECALL), and Multi-Object Tracking Accuracy (MOTA). Additionally, we verify the visual fidelity of our synthesized samples using both the Fréchet Inception Distance (FID) [14] and the Fréchet Video Distance (FVD) [39].
4.2 Implementation Details
SubjectDrive employs a two-stage video generation strategy: the first stage generates images, and the second stage generates videos. The first stage is optimized for 56k steps, while the second stage is optimized for 84k steps. During sampling, we utilize the Denoising Diffusion Implicit Models (DDIM) sampler, configured to 25 steps, to yield video clips at a resolution of $256\times 512$ across 8 frames. Furthermore, we have chosen StreamPETR with a ResNet50 [13] backbone as our evaluating perception model, which is trained at a resolution of $256\times 512$.
Table 1: Evaluation of data scaling on detection and tracking tasks.
#Extra Samples    Source           Tracking         Detection
@@ 4.2 Implementation Details
0 nuScenes [5] 30.1 1.379 46.9 34.5
93k nuImages [5] 33.8 (+3.7%) 1.324 49.5 (+2.6%) 37.8
23k Panacea[43] 33.7 (+3.6%) 1.353 49.2 (+2.3%) 37.1

### v2

# SubjectDrive: Scaling Generative Data in Autonomous Driving via Subject Control

@@ Abstract
Introduction
@@
Deep generative models (Rombach et al. 2022; Blattmann et al. 2023b; Yu et al. 2023; Zhang, Rao, and Agrawala 2023; Song, Meng, and Ermon 2020; Ho, Jain, and Abbeel 2020) have achieved remarkable progress recently, demonstrating excellence in generating high-quality and realistic visual content. Diffusion models(Song, Meng, and Ermon 2020; Ho, Jain, and Abbeel 2020), a key contributor to this advancement, are renowned for their stable and top-quality sample generation. Through the utilization of modern diffusion models, the recent breakthroughs in controllable technology (Zhang, Rao, and Agrawala 2023) now facilitate precise and flexible content customization. These developments enable the creation of synthetic samples that are almost indistinguishable from real, human-annotated data. By leveraging the generative data, methods dependent on supervised training can be substantially enhanced for tasks such as image classification(Fan et al. 2024), object detection(Wang et al. 2024a), and object tracking(Li et al. 2023a). Motivated by these successes, there is a growing interest in applying generative models to the synthesis of natural data for more discriminative tasks (Sarıyıldız et al. 2023; Azizi et al. 2023; Nguyen et al. 2024).
Figure 1: The comparison between existing method and SubjectDrive framework.(a) Existing data generation framework that uses the control sequence and sampling noise to generate synthetic data with limited sample diversity and scalability. (b) Compared with the traditional framework, SubjectDrive introduces additional synthesis diversity by incorporating extra subject control to enhance the scalability of generative model. (c) Evaluation of data scaling on the nuScenes detection and tracking task.
Autonomous driving, as a key application domain, heavily relies on large-scale annotated data and has increasingly paid attention to generative data collection methods. One of the most challenging tasks in autonomous driving is BEV (Bird’s Eye View) perception, which includes 3D object detection(Wang et al. 2023a; Pang et al. 2023; Li et al. 2022; Liu et al. 2022, 2023a), 3D object tracking(Pang et al. 2023; Fischer et al. 2022; Marinello, Proesmans, and Van Gool 2022), and 3D lane detection. Acquiring annotated training data for these tasks is not only costly but also burdened by significant concerns regarding data privacy and usage rights, creating barriers that hinder data collection, labeling, release, and exchange processes, ultimately slowing progress in the field. To mitigate this issue, previous work has developed generative solutions such as BEVGen (Swerdlow, Xu, and Zhou 2023) and BEVControl (Yang et al. 2023), which could produce annotated training samples by adapting an image synthesis model. More recent efforts (Gao et al. 2023; Wen et al. 2023; Li, Zhang, and Ye 2023; Lu et al. 2023) have ventured into crafting driving scene videos, which successfully bolster more sophisticated temporal BEV perception models (Wang et al. 2023a).
Albeit promising, most existing work focuses on exploring the potential of generative data on a small scale, which fails to fully exploit the capacity of generative models to produce an inexhaustible supply of training samples. Notably, although early methods have demonstrated superior performance when incorporating generated samples alongside real data (Wen et al. 2023), there is still a gap compared to using existing large-scale real image data (Caesar et al. 2020) for pre-training, as illustrated in Fig. 1(a).
To further ignite the potential of synthetic data, this paper will first investigate the scalability of generative data produced by existing methods. As shown in Fig. 1(a), we observe that when the number of synthetic samples generated by the existing method increased from 46K to 69K, the performance in 3D object tracking notably decreased. We found that this lack of scalability is intrinsically related to the limited diversity of the generated samples. As shown in Fig. 2, different generative samples with the same annotation exhibit remarkably similar appearances in their foreground objects, i.e., the car, despite differences in the background. We also found that when the diversity issue is addressed by the method we introduce later, the scalability of generative data can be significantly improved.
@@
In this paper, we tackle the challenge of scaling generative data for autonomous driving. We present SubjectDrive, an advanced video generation framework designed to enhance the scalability of generative models. Our initial findings reveal that conventional video generation pipelines struggle to scale effectively with increased data volumes due to its limited diversity. To overcome this limitation, we propose a novel generation framework centered on augmenting sampling diversity. Specifically, we integrate a feature termed subject control into existing generation pipelines. This feature empowers generative models to manipulate the diversity of the synthesis process by providing a mechanism to dictate the visual appearance of foreground elements in generated samples. This innovative feature enables the blending of the inherent stochastic nature of the generative process with the diversity drawn from external data sources, thereby crafting a more powerful model capable of producing scalable and diverse samples.
Figure 2: Above: Existing autonomous driving generative models struggle to produce diverse foreground samples. Below: By enhancing the sampling diversity capabilities with our subject control methods, the diversity of the generated foreground sample has significantly improved.
@@
More concretely, SubjectDrive is designed around a trio of innovative modules that collectively enable robust subject control capabilities. Initially, the model leverages a Subject Prompt Adapter, integrating subject control with the existing text-conditioned branch. Subsequently, to enhance the model’s ability to inject spatial information, we introduce a Subject Visual Adapter that directly utilizes visual features, incorporating them into the existing diffusion U-Net architecture. Lastly, to ensure consistent injection of these features over time, we deploy Augmented Temporal Attention that expands the model’s temporal-spatial context. Together, these modules empower SubjectDrive to perform subject-conditioned video generation with compelling efficacy. Overall, our research includes three main contributions:
  • Unlike existing synthetic methods for autonomous driving, which are limited to small-scale validations, we propose a novel and critical research direction: investigating the scalability of generative data in autonomous driving. Our exploration provides a deeper understanding of the challenges and solutions associated with leveraging large-scale synthetic data in this domain.
  • To address the challenges faced by existing methods during data scaling, particularly the issue of limited diversity, we introduce a novel generative framework called SubjectDrive. This framework enhances sampling diversity and improves the scalability of generative data through an innovative mechanism known as subject control.
@@
  • Extensive experiments on the nuScenes dataset (Caesar et al. 2020) validate the effectiveness of our proposed method. Remarkably, our approach is the first generative technique to surpass the performance of perception models pre-trained on a large-scale real dataset, specifically nuImages(Caesar et al. 2020). These outstanding results underscore the transformative potential of generative data in advancing autonomous driving technologies, offering a promising direction for future research in this field.
Related Work
Scalable Data Synthesis for Autonomous Driving
@@
The use of synthetic data has been widely explored for visual perception tasks that require great of efforts in labeled data collection. Examples include applications in image classification (Azizi et al. 2023; Sarıyıldız et al. 2023), semantic segmentation (Nguyen et al. 2024), and object tracking (Li et al. 2023b). Recently, there has been a burgeoning interest in employing generative data for the challenging BEV perception tasks (Wang et al. 2023a; Li et al. 2022), which demand precise geometric and appearance alignment from generative models. Initial works aim at synthesizing street-view images (Swerdlow, Xu, and Zhou 2023; Yang et al. 2023), with subsequent studies extending into the generation of driving scene videos (Wen et al. 2023; Lu et al. 2023; Wang et al. 2023b). Despite these successful efforts, the majority of existing research has been limited to producing a small quantity of training samples, not fully exploiting the generative models’ capacity to offer an unlimited reservoir of samples. Data scaling efforts (Azizi et al. 2023) are sparse and, thus far, not particularly successful. It’s imperative to recognize that recent breakthroughs highlight the paramount importance of data scaling in training sophisticated deep learning models, such as GPT (Achiam et al. 2023), SVD (Blattmann et al. 2023a), and Sora (Brooks et al. 2024). Therefore, addressing the challenge of amplifying generative data volumes is of essence. This work delves into scalable data synthesis for autonomous driving, providing analysis results and proposing an enhanced framework.
Diffusion-based Generative Models with Subject Controls
@@
Diffusion models with subject control are designed to embed target subjects into generated visual content, guided by reference images (Choi et al. 2023). These approaches (Han et al. 2023; Kumari et al. 2023; Li et al. 2023c; Chen et al. 2023b; Ma et al. 2023; Xiao et al. 2024) initially emerged in image generation research to facilitate identity-preserving applications. Early methods (Han et al. 2023; Kumari et al. 2023; Li et al. 2023c; Gal et al. 2022; Ruiz et al. 2023) adopted a fine-tuning framework to adjust a specific model capable of generating images of the desired subject. A notable example is DreamBooth (Ruiz et al. 2023), which fine-tunes a diffusion model using a small set of subject images. While these fine-tuning methods achieved high-quality results, they were limited by the significant resources required for model tuning, making them impractical for scenarios with numerous target entities. To address these limitations, tuning-free methods (Ma et al. 2023; Ye et al. 2023) were developed. Subject Diffusion (Ma et al. 2023), for instance, achieved impressive customization by introducing a subject conditioning module and training on a bespoke dataset of subject pairs. Subsequent works, such as Cones2 (Liu et al. 2023b) and others (Chen et al. 2023b), have expanded tuning-free subject control to scenarios involving multiple subjects.
@@
In the video domain, there is also growing interest in subject-controllable generation (Jiang et al. 2023; Chen et al. 2023a; Wang et al. 2024b). For instance, VideoBooth (Jiang et al. 2023) proposed a framework capable of generating consistent videos containing the subjects specified in image prompts. Additionally, CustomVideo (Wang et al. 2024b) introduced a multi-subject-driven text-to-video model powered by a simple yet effective co-occurrence and attention control mechanism. Our work falls into the category of subject-driven video generation. However, unlike previous efforts, we explore the significance of such subject-controlled generation for data scaling, especially in the context of autonomous driving.
Method
Figure 3: Overview of SubjectDrive. The pipeline of SubjectDrive involves a frozen auto-encoder and a trainable UNet-based diffusion model. Different control signal sources including extended text prompt, condition layout, and subject bank.
Preliminary: Latent Diffusion Models
@@
Diffusion Models (DMs) (Song, Meng, and Ermon 2020; Ho, Jain, and Abbeel 2020) iteratively denoise a random noise with Gaussion distribution $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ over T steps to generate target data according to Eq. 1, where the functions $\mu_{\theta}$ and $\Sigma_{\theta}$ are derived from the denoising model $\epsilon_{\theta}$. The training of DMs includes both a diffusion process and a denoising process. The diffusion process begins by adding Gaussion noise to the original data sample $x_{0}$ over $t$ steps, controlled by the scheduled noise strength $\beta_{t}$, which can be simplified as Eq. 2 where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\left(1-\beta_{t}\right)$. To learn the denoising process, the denoising model $\mathbf{\epsilon}_{\theta}$, which aims to estimate the original noise $\mathbf{\epsilon}$ from the noisy data $x_{t}$, is optimized by minimizing the loss function as shown in Eq. 3.
@@
Given the significant computational burden of diffusion models in generating high-resolution images or videos, Latent Diffusion Models (LDMs) (Rombach et al. 2022) have become popular in recent studies. With a pre-trained auto-encoder (Kingma and Welling 2013) to handle the compression and reconstruction between high-dimensional visual data and low-redundancy latent representations, the diffusion model can concentrate exclusively on generation in latent space, significantly reducing computational costs. Consequently, our method employs LDMs for video generation.
SubjectDrive
SubjectDrive is designed to enhance the scalability of generative data, thereby promoting perception models for autonomous driving applications. Despite the ability of advanced generative methods (Wen et al. 2023) to produce high-quality driving-scene videos, they narrowly uplifts the performance on downstream perception tasks as illustrated in Fig. 1(c). We believe this is primarily due to the limited diversity of generated foreground elements, which are crucial for autonomous driving. Thus, different from tradition generation pipeline using control sequence to guide global scene generation, we innovatively integrate subject control mechanism into generation process, allowing for the injection of external subjects from extensive open-source data. The integration of controlled subjects not only boosts controllability but also effectively enhances the diversity of generated foreground elements, which shows strong augmentation for autonomous driving applications.
@@
Overview
Our framework is built on a text-to-video diffusion model, i.e. Panacea (Wen et al. 2023), which is a strong baseline for multi-view video generation. The overall training framework of SubjectDrive is illustrated in Fig. 3. The diffused noisy input is fed into the trainable diffusion model to generate latent video under the guidance of text, condition layout, and injected subjects. We inherit the guidance of condition layout with ControlNet (Zhang, Rao, and Agrawala 2023) from Panacea. For subjects guidance, we first extend text prompt and augment the text embedding of subject part with Subject Prompt Adapter. On the other hand, we insert the gated self-attention layer into diffusion model to enhance the location guidance of subjects captured by Subject Visual Adapter. Furthermore, we replace the conventional temporal 1D attention with proposed Augmented Temporal Attention to improve the temporal consistency of injected subjects.
Subject Prompt Adapter
@@
Figure 4: The Subject Prompt Adaptor which augments the original text embedding of extended prompt with corresponding ID identifier and visual semantic information to enhance the expressity of subjects.
To achieve this, we first extend the scene prompt with injected subject category description. For example, the prompt for the $N_{th}$ frame that contains $M$ subjects is organised as ”[Scene Prompt], including [Subject $X_{1}$], [Subject $X_{2}$]…. [Subject $X_{M}$].[Subject $X_{i}$] here refers to the category word of subject $X_{i}$, e.g., car, bus. The extended prompt is input into the pre-trained CLIP (Radford et al. 2021) text encoder to obtain the original text embedding $z^{t}\in{\mathbb{R}^{L\times d}}$, where $L$ represents the length of the text embedding and $d$ is the dimension of the text embedding. To further integrate subject attributes for better expressity, we extract the ID identifier of each subject from condition layout to ensure coherent trajectories across frames, and encode it with a learnable codebook and an ID-adapter comprised of MLP to get the ID embedding $z_{i}^{id}\in{\mathbb{R}^{d}}$ for subject $X_{i}$. Similarly, we encode the image of $X_{i}$ by the CLIP image encoder and an image-adapter to obtain the corresponding visual semantic embeddings $z_{i}^{v}\in{\mathbb{R}^{d}}$.
@@
$\hat{z}_{i}^{t}=\mathrm{MLP}\left([z_{i}^{t}+z_{i}^{id},z_{i}^{v}]\right),\ \ i\in\left\{\mathrm{Index}_{1,2,\cdots,M}\right\}$ (4)
@@
Subject Visual Adapter
@@
Figure 5: The Subject Visual Adapter which injects location-enhanced subject information into visual features cooperated with gated self-attention.
Inspired by (Li et al. 2023d), SVA first combines subject images and corresponding locations to form location-enhanced subject embeddings, and then injects them into frames with a control signal to guide the generation of subjects at specified location. To obtain the location-enhanced subject embedding $f^{vl}$, the location of each subject, represented by its coordinates $l=[x_{min},y_{min},x_{max},y_{max}]$ in the current frame, is encoded with fourier embedding, further integrated into visual embedding $f^{v}$ of corresponding subject as
$f^{vl}=\mathrm{MLP}[f^{v},\mathrm{Fourier}(l)]).$ (5)
To inject the location-enhanced subject embedding into generated frame, we employ the gated self-attention layer (Li et al. 2023d) that locates between each paired self-attention layer and cross-attention layer in the UNet. In the gated self-attention layer, the location-enhanced subject embedding $f^{vl}$ interacts with frame visual tokens $z$ by an attention operation to capture the dependency between subject embedding and visual tokens, followed by a token selection operation $TS$ to preserve only visual tokens. To adaptively adjust the guidance scale of location information over frames, a gating factor is learned which is operated as
$z=z+\tanh(\gamma)\cdot\mathrm{TS}(\mathrm{SelfAttn}([z,f^{vl}])),$ (6)
@@
Augmented Temporal Attention
The Augmented Temporal Attention is designed to capture long-range movement of subjects with feasible computation cost.
Figure 6: The Augmented Temporal Attention integrates conventional temporal 1D attention with decomposed attention on temporal-horizontal (TX) plane and temporal-vertical (TY) plane to effectively capture long-range movements of subjects.
@@
$z^{o}=\mathrm{SelfAttn}(z^{TX})+\mathrm{SelfAttn}(z^{TY})+z^{T}+z^{i}.$ (7)
Experiment
Evaluation Datasets and Metrics
@@
Evaluation Metrics. Following Panacea, we evaluate two key objectives using perceptual models. First, we assess the generative model’s controllability by measuring how well the generated data aligns with BEV annotations, using a pre-trained StreamPETR model to compare performance against real data. Second, we evaluate how the generated data improves perceptual model performance on detection and tracking tasks. Detection metrics include nuScenes Detection Score (NDS), mean Average Precision (mAP), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Tracking metrics include Average Multi-Object Tracking Accuracy (AMOTA), Precision (AMOTP), Recall, and Accuracy (MOTA). Visual fidelity is validated with Fréchet Inception Distance (FID)(Heusel et al. 2017) and Fréchet Video Distance (FVD)(Unterthiner et al. 2018).
Implementation Details
SubjectDrive adopts a two-stage video generation approach: image generation in the first stage (optimized for 56k steps) and video generation in the second (84k steps). Experiments are conducted on 8A100 GPUs using the DDIM sampler with 25 steps to produce $256\times 512$ resolution video clips spanning 8 frames. The evaluation uses StreamPETR with a ResNet50 backbone (He et al. 2016), trained at $256\times 512$ resolution. To construct the subject bank, we draw from two distinct sources, categorizing them into internal and external subject banks. The internal subject bank is curated by collecting subjects from the training set of the nuScenes dataset. The external subject bank is established by integrating external vehicle datasets from the open-source CompCars (Yang et al. 2015) dataset. The internal subject bank is used during the training phase, while the external subject bank is mixed with the internal one during the sampling phase to promote the generation of diverse data.
Main Results
Analysis of Synthetic Data Scale-up. We investigate the impact of scaling up the quantity of generative data on the performance of downstream tasks and present the results in Table 1. Two interesting findings emerge. First, increasing the quantity of synthetic data has a positive influence on the performance of the perception model. Although using a small amount of synthetic data is initially less effective than using real nuImages data, increasing the data volume can boost the models to achieve superior performance in both NDS and AMOTA. This demonstrates that scaling up generative data is essential to unlock the potential of generative data production. Second, the integration of external subjects can substantially improve scaling performance. For instance, while tripling the generative data from Panacea yields a mere 0.2 improvement in NDS, incorporating SubjectDrive at the same data scale elevates the NDS by 1.0. Similar trends are also observed in AMOTA. These results demonstrate that utilizing external subjects is an effective way to enhance the generative model’s scaling capability.
#Extras    Source          Tracking        Detection
@@
0 nuScenes 30.1 1.379 46.9 34.5
93K nuImages 33.8 (+3.7%) 1.324 49.5 (+2.6%) 37.8
23K Panacea 33.7 (+3.6%) 1.353 49.2 (+2.3%) 37.1